# A Fine-grained Sentiment Analysis of App Reviews using Large Language Models: An Evaluation Study

### v2

@@ Abstract
Analyzing user reviews for sentiment towards app features can provide valuable insights into users’ perceptions of app functionality and their evolving needs. Given the volume of user reviews received daily, an automated mechanism to generate feature-level sentiment summaries of user reviews is needed. Recent advances in Large Language Models (LLMs) such as ChatGPT have shown impressive performance on several new tasks without updating the model’s parameters i.e. using zero or a few labeled examples. Despite these advancements, LLMs’ capabilities to perform feature-specific sentiment analysis of user reviews remain unexplored. This study compares the performance of state-of-the-art LLMs, including GPT-4, ChatGPT, and LLama-2-chat variants, for extracting app features and associated sentiments under 0-shot, 1-shot, and 5-shot scenarios. Results indicate the best-performing GPT-4 model outperforms rule-based approaches by 23.6% in f1-score with zero-shot feature extraction; 5-shot further improving it by 6%. GPT-4 achieves a 74% f1-score for predicting positive sentiment towards correctly predicted app features, with 5-shot enhancing it by 7%. Our study suggests that LLM models are promising for generating feature-specific sentiment summaries of user reviews111https://github.com/Faiz-UT/Eval-Feature-Sentiment-Extraction-LLMs.
@@
RQ1: How effective are Large Language Models (LLMs) at extracting feature-sentiment pairs from user reviews under zero-shot settings?
@@
RQ2: How effective are Large Language Models (LLMs) at extracting feature-sentiment pairs from user reviews under few-shot settings?
@@ 4.3 Baselines
We compared the performance of our LLM-based models with four state-of-the-art baselines that used the same evaluation procedures, such as labeled dataset and feature matching strategies, as our study. Each baseline approach is briefly described as follows:
@@ 5 Results
This section answers our RQs by presenting the performance results of LLMs for extracting app features and corresponding sentiments from app reviews under zero-shot (i.e. RQ1) and few-shot (i.e. RQ2) settings.
5.1 RQ1 - Zero-shot performance of LLMs at extracting feature-sentiment pairs
@@ 5.1.2 Feature-specific sentiment prediction performance
5.2 RQ2 - Few-shot performance of LLMs at extracting feature-sentiment pairs
This section presents the few-shot (i.e. 1-shot and 5-shot) performances of LLMs, including feature-extraction and feature-specific sentiment prediction results.
@@ 6.2 Implications for Practitioners
With the vast landscape of both open-source and proprietary LLM models, selecting the most suitable LLM for integration into tools designed for fine-grained analysis of user feedback can be challenging for practitioners. Our evaluation of state-of-the-art (SOTA) LLMs offers a realistic assessment of their performance. Furthermore, the varying performance of LLM models across different applications, as illustrated in Figure 4, underscores potential generalization issues with LLMs.
@@ Acknowledgment
This work has received funding from the EU H2020 program under the SoBigData++ project (grant agreement No. 871042), by the CHIST-ERA grant No. CHIST-ERA-19-XAI-010, (ETAg grant No. SLTAT21096), and partially funded by HAMISON project.
@@
| Identify all feature expressions with their corresponding sentiment polarities in | |
| the following review text (enclosed in double quotations). Output the results in the | |
| format of [(’feature expression’,’sentiment polarity’), …]. If no feature expression | |

### v3

# How Effectively Do LLMs Extract Feature-Sentiment Pairs from App Reviews?

@@ Abstract
Automatic analysis of user reviews to understand user sentiments toward app functionality (i.e. app features) helps align development efforts with user expectations and needs. Recent advances in Large Language Models (LLMs) such as ChatGPT have shown impressive performance on several new tasks without updating the model’s parameters i.e. using zero or a few labeled examples, but the capabilities of LLMs are yet unexplored for feature-specific sentiment analysis. The goal of our study is to explore the capabilities of LLMs to perform feature-specific sentiment analysis of user reviews. This study compares the performance of state-of-the-art LLMs, including GPT-4, ChatGPT, and different variants of Llama-2 chat, against previous approaches for extracting app features and associated sentiments in zero-shot, 1-shot, and 5-shot scenarios. The results indicate that GPT-4 outperforms the rule-based SAFE by 17% in f1-score for extracting app features in the zero-shot scenario, with 5-shot further improving it by 6%. However, the fine-tuned RE-BERT exceeds GPT-4 by 6% in f1-score. For predicting positive and neutral sentiments, GPT-4 achieves f1-scores of 76% and 45% in the zero-shot setting, which improve by 7% and 23% in the 5-shot setting, respectively. Our study conducts a thorough evaluation of both proprietary and open-source LLMs to provide an objective assessment of their performance in extracting feature-sentiment pairs. 111https://github.com/Faiz-UT/Eval-Feature-Sentiment-Extraction-LLMs.
@@
RQ1: How does the zero-shot performance of LLMs compare to existing methods for extracting feature-sentiment pairs from app reviews?
@@
RQ2: How does the few-shot performance of LLMs compare to zero-shot and existing methods for extracting feature-sentiment pairs from app reviews?
@@ 4.3 Baselines
We compared the performance of our LLM-based models with four state-of-the-art baseline methods briefly described in this section. Since Dąbrowski et al.’s study [25] evaluated these baseline methods using the same labeled dataset and evaluation procedure as our study, we used their performance results rather than re-running the baseline methods ourselves.
@@ 5 Results
This section answers our RQs by comparing the zero-shot and few-shot performance results of LLMs with baseline methods for extracting app features and corresponding sentiments from app reviews.
5.1 RQ1 - Comparison of zero-shot LLM performance and baseline methods for extracting feature-sentiment pairs
@@ 5.1.1 Feature extraction performance
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: GPT-4 surpasses SAFE by 17% in f1-score. However, the fine-tuned RE-BERT outperforms GPT-4 by 12% in f1-score.
@@ 5.1.2 Feature-specific sentiment prediction performance
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: To predict positive and neutral sentiments, GPT-4 achieves the best f1 scores of 76.1% and 44.8%, respectively; while LLama-2-70B yields the best f1-score of 50.4% for negative sentiment prediction.
5.2 RQ2 - Comparison of few-shot LLM performance against zero-shot and baseline methods for extracting feature-sentiment pairs
We compare the few-shot (i.e. 1-shot and 5-shot) performance of LLMs with zero-shot and baseline methods for extracting feature-sentiment pairs.
@@ 5.2.1 Feature extraction performance
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: With 5-shot learning, GPT-4 improved the f1-score by 6% (i.e., 56%), representing a 23% improvement over the SAFE approach. However, the fine-tuned RE-BERT still outperforms GPT-4 by 6% in f1-score.
@@ 5.2.2 Feature-specific sentiment prediction performance
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: For neutral sentiment prediction, 5-shot improves the f1-score by 23% for GPT-4 and 14% for LLama-2-70B. In positive sentiment prediction, 5-shot improves the f1-score by 7% for GPT-4 and 3% for LLama-70b. However, for negative sentiment, 5-shot does not improve the performance of GPT-4, and f1-score decreases for ChatGPT and LLama-2-70B by 7% and 8%, respectively.
@@ 6.2 Implications for Practitioners
Our findings indicate that while LLMs show promise, their precision and recall in extracting app features from user reviews are not yet adequate for practical implementation. Although prompt engineering demonstrates lower performance compared to fine-tuned pretrained models, it offers a more cost-efficient approach for practitioners, particularly in settings where labeled data is scarce. Using LLMs for prompting is highly suited to extracting feature-sentiment data from a range of feedback sources, including in-app feedback, community forums, emails, and Reddit, which all demand a broad understanding of language.
@@ Acknowledgment
This work has received funding from the EU H2020 program under the SoBigData++ project (grant agreement No. 871042), CHIST-ERA grant No. CHIST-ERA-19-XAI-010, grant PRG1226 of the Estonian Research Council, and the HAMISON project.
@@
| As an expert information extractor, identify all feature expressions with their corresponding | |
| sentiment polarities in the following review text (enclosed in double quotations). Output the | |
| results in the format of [(’feature expression’,’sentiment polarity’), …]. If no feature expression | |